You have an absolutely brilliant child who is six years old and will be attending a private college in twelve years. You know that a four-year college now costs at least $30,00 per year, including tuition, books, and room and board. The cost of sending a child to college has increased by 7 percent per year, and you believe this will be true for the next twelve years. How much will the annual tuition be when your child is eighteen?

a. $67,565.75

b. $66,320.44

c. $60,365.89

d. $45,000
Answer #1

A = P(1+r/100)^n            

A = 30,000*(1+7/100)^12          

A = 30,000*(1.07)^12

A = 30,000*2.2522       

A = 67,565.75

The answer is a.